# Exploiting Meta-Learning-based Poisoning Attacks for Graph Link Prediction

### v1

@@ 6       Compute attack loss $\widetilde{L}_{i}$;
7       $\widetilde{L}$ = $\widetilde{L}$ + $\widetilde{L}_{i}$;
@@ 8       Back propagation on $\nabla$: update $\theta_{i}\rightarrow\theta_{i+1}$;
9
10Compute the meta-gradient $\widetilde{\nabla}$ with respect to attack target based on $\widetilde{L}$ ;
11 $A_{adv}=$ Adversarial modification(~$\widetilde{\nabla},A,\Delta$);
12
13return $A_{adv}$;
@@ 1 Initial $A_{adv}=A$
2 Symmetry the gradient matrix $\widetilde{\nabla}$ = ($\widetilde{\nabla}$ + ~$\widetilde{\nabla}^{T}$)/2
@@
4       Get the index of the largest magnitude in $\widetilde{\nabla}$:$(i,j)$ = $|\widetilde{\nabla}|$);
5       if ~$\widetilde{\nabla}_{ij}>0$ and $A_{ij}=0$ then
@@ 7             $\Delta$ = $\Delta$ 1;
9      if ~$\widetilde{\nabla}_{ij}<0$ and $A_{ij}=1$ then
@@ 4.4 Attack Strategy
The primary purpose of the attack model is to extract information by mimicking the original process since the attacker lacks access to the original model and the training process. The essence of the attack model (Algorithm 1) lies in leveraging meta-learning during its training phase, enabling the attacker to gain a comprehensive understanding of the process. The attack model is trained on the original graph data using weighted cross-entropy (Algorithm 1, lines 3 and 4), defined as:
@@
The attack loss is aggregated as meta-learning approach during each epoch (Algorithm 1, line 6). The key part of the attack model is that the attacker considers each training epoch as a task (episode) and intends to learn from all tasks (epoches). Instead of looking into a specific phase in the training process, meta-learning helps the attacker to capture the overall information of the whole training. The attacker tries to maximize the attack loss with respect to the graph structure ($Adj$) to conduct the effective attack on the graph structure. At the end of the training, the attacker computes the meta-gradient from the attack loss for further processing(Algorithm 1, line 8).
@@
Table 1: Results of the Targeted Attack for Link Adversarial Dissolution.
Cora Citeseer Pubmed
Attack Budget
1% 2.5% 5% 1% 2.5% 5% 1% 2.5% 5%
Attack Approach ASR
Null Model [23] 10% 10% 10% 10% 10% 15% 10% 10% 15%
IGA [14] 25% 30% 30% 30% 30% 35% 25% 25% 30%
PGD [12] 35% 35% 40% 40% 40% 40% 35% 35% 40%
MetaBase [16] 40% 40% 45% 40% 40% 40% 35% 35% 45%
DICE [13] 30% 35% 40% 35% 35% 40% 35% 35% 35%
Our approach 45% 50% 50% 45% 45% 50% 40% 45% 50%
Table 2: Results of the Targeted Attack for Link Adversarial Activation.
Cora Citeseer Pubmed
Attack Budget
1% 2.5% 5% 1% 2.5% 5% 1% 2.5% 5%
Attack Approach ASR
Null Model [23] 15% 20% 20% 15% 15% 20% 15% 15% 15%
IGA [14] 55% 55% 60% 50% 60% 60% 45% 55% 60%
PGD [12] 60% 60% 60% 60% 60% 65% 55% 60% 60%
MetaBase [16] 65% 70% 70% 60% 65% 70% 65% 65% 65%
DICE [13] 60% 65% 65% 50% 60% 65% 55% 55% 60%
Our approach 100% 100% 100% 100% 100% 100% 90% 90% 90%
5 Experimental Results
@@
5.1 Experimental Setup
To evaluate the effectiveness and performance of our proposed attack method, we implemented the entire architecture using PyTorch. Our experiments aimed to answer the following questions:
  • Does our approach reduce the link prediction accuracy of VGAE?
  • Is our approach effective in various attack scenarios (objectives)?
@@ 5.1 Experimental Setup
To answer these questions, we deployed a two-layer VGAE to evaluate the effectiveness of our poisoning attack approach on three widely used graph datasets. By training VGAE separately on the original and poisoned graph data, we can observe any disparities in link prediction performance, indicating the effectiveness of our attack. In addition, to assess adaptability and robustness, we conducted evaluations under two distinct attack scenarios, each aligned with a specific attack objective. We compared the efficacy of our approach with six state-of-the-art methods, namely PGD [12], DICE [13], IGA [14], CLGA [17], MCGCL [24] and MetaBase [16]. Since CLGA [17] and MCGCL [24] are both designed based on the embedding of the graph that inherently considers the entire graph. It is challenging to isolate the single link as a target in their framework. Therefore, we limit our comparison of these two methods to untargeted attacks.
In our experiment, the attacker also utilizes a VGAE as the attack model to apply the adversarial modification on graph data and maximize the ability of our attack. As mentioned in the previous section, the adversarial modification is constrained by the attack budget, and we investigate the attack budget as a key parameter in our experiments. To ensure that the attack remains unnoticeable, we set the attack budget within a small range, varying from 1% to 5% of the total links in the dataset to produce the poisoned data for our poisoned attack. For each experiment, we conducted original training using original data and poisoning training using poisoned data. We then test both models with the same test data to evaluate their link prediction performance. The performance of the original model will be used as a baseline. The difference in performance between these two models indicates the effectiveness of our proposed attack. Moreover, we also leverage the null model [23] as an additional baseline. In the null model, the total number of links in the graph is fixed, and adversarial graphs are generated by randomly removing and reconnecting links within the original graph. Comparing our attack approach with the null model allows us to evaluate whether our approach explicitly highlights the significant vulnerable points in the graph structure.
Figure 3: Different adversarial attack scenarios under targeted attack.
5.2 Experimental Dataset
In our experiments, we utilized three widely used graph datasets sourced from public repositories to evaluate the performance of our proposed attack approach. These datasets represent scientific publications, each node representing a publication, and each link denoting citations.
@@ 5.2 Experimental Dataset
  • Cora: The Cora dataset [25] contains scientific publications that are categorized into 7 classes. It has 2708 nodes and 5429 links. Each node is described by a word vector that indicates the presence of certain words in the publication from a predefined dictionary.
  • Citeseer: The Citeseer dataset [26] is a collection of scientific publications classified into 6 classes. It has 3312 nodes and 4732 links. Each node is described by a word vector that indicates the presence of certain words in the publication from a predefined dictionary.
  • Pubmed: The Pubmed dataset [27] comprises scientific publications from the PubMed database categorized into 3 classes. It has 19717 nodes and 44338 links. Each node is described by a Term Frequency-Inverse Document Frequency (TF-IDF) weighted word vector to capture the importance of words in the publication.
@@
(b) Citeseer dataset
(c) Pubmed dataset
Figure 4: ROC-AUC Results of the Untargeted Attack Scenario.
@@
(b) Citeseer dataset
(c) Pubmed dataset
Figure 5: AP Results of the Untargeted Attack Scenario.
5.3 Attack Scenarios
As elaborated in Section 3, adversarial attacks can manifest in two distinct scenarios (based on the objectives of the attack): targeted attacks and untargeted attacks. In our experiments, we address both scenarios as follows:
5.3.1 Targeted Attack
In this scenario, the attacker aims to manipulate the prediction of a single link within the graph, which can be categorized into two types as depicted in Fig. 3:
  • Link adversarial dissolution: In this attack, the objective is to alter/manipulate the prediction result of a link from present (1 - link exists) to absent (0 - link does not exist) in the adjacency matrix, thereby affecting the targeted position.
  • Link adversarial activation: In contrast, this attack seeks to change the prediction of a link from absent (0 - link does not exist) to present (1 - link exists) in the adjacency matrix.
Generally, when observing the performance of targeted attacks, most works focus on overall attack performance. However, it is crucial to investigate both link activation and dissolution caused by our attacks because these phenomena can have vastly different implications across various applications. For example, in social networks, the activation of links can foster stronger relationship ties, enhancing community engagement and collaboration. In contrast, the dissolution of links might isolate relationships, leading to fragmentation and decreased user interaction. Therefore, we evaluate both cases separately to help us better understand and address the potential vulnerabilities of link prediction.
5.3.2 Untargeted Attack
In this scenario, the attacker focuses on a specific set of links (e.g., testing set) of the entire graph, aiming to maximize the misclassification rate within that set.
5.4 Evaluation Metric
  • Attack Success Rate (ASR): For targeted attacks, ASR indicates the proportion of successful targeted attacks out of all attempted targeted attacks. This metric reflects the overall attack strength of the proposed attack, as shown in Eq. (4).
    $ASR=\frac{all\>success\>attack\>attampts}{all\>attack\>attampts}$ (4)
@@ 5.4 Evaluation Metric
  • ROC-AUC and AP Score: In untargeted attacks, we utilize the Receiver Operating Characteristic (ROC), Area Under the Curve (AUC), and Average Precision (AP) Score to evaluate the performance of the model after the poisoning attack. ROC-AUC score provides a comprehensive understanding of the model’s ability to distinguish between positive and negative links. The ROC-AUC score is threshold-independent, evaluating model performance across all possible thresholds, and is more resilient to class imbalance compared to accuracy, making it a robust measure. Additionally, it provides a clear basis for comparing multiple models. Moreover, as graph data are usually considered imbalance data, the AP score emphasizes the precision-recall trade-off of the model. By comparing the original model’s performance with its post-attack performance, the effectiveness of the poisoning attack is clearly demonstrated, highlighting its impact on the robustness and performance of the model.
5.5 Effectiveness Analysis
5.5.1 Evaluation of targeted attack
Tables 1 and 2 present the success rates of 20 attack attempts. All the state-of-the-arts and our approach outperform the null model baseline. This shows that all these attack approaches identify and exploit vulnerable parts. A higher ASR indicates a stronger targeted attack ability on individual target links in the link prediction task. Our approach demonstrates remarkable effectiveness in the targeted attack scenario, outperforming other state-of-the-art methods. Additionally, our approach achieves the attacker’s objective with a smaller attack budget, making it more inconspicuous. As the attack budget increases in targeted attacks, there is a slight increment in the targeted attack results. This could be attributed to the fact that for a targeted attack on a single link, fewer modifications are needed to degrade the prediction accuracy of that link.
Another notable observation from the targeted attack results is that the ASR for link adversarial dissolution is significantly higher than that for link adversarial activation. This suggests that breaking an existing link is more challenging than creating a new adversarial link, both in our attack and in other state-of-the-art methods. This intuitively aligns with real-world scenarios; forming a new connection between two entities is generally simpler than breaking an existing relationship. For instance, in a social network, it is generally simpler to establish a new connection between two individuals than to dissolve an existing relationship between them.
5.5.2 Evaluation of untargeted attack
The results of the untargeted attack in the three datasets are depicted in Fig. 4 and Fig. 5. In the experiments, we used two baselines, the original model and the null model. The original model represents the performance of the link prediction on the original graph data. It is unaffected by the attack budget as it is clean data. The effectiveness of the attack methods is evaluated by observing the drop in ROC-AUC and AP scores. The null model serves as a benchmark to evaluate whether the approach is capable of highlighting the vulnerabilities of the graph data. All approaches demonstrate their effectiveness compared to the original model and null model. Our method shows superior performance in untargeted attack scenarios, slightly surpassing other state-of-the-art approaches. In particular, our method achieves a lower ROC-AUC and AP score while requiring a smaller attack budget. Unlike targeted attacks, the impact of the attack budget is more pronounced in untargeted attacks. As the attack budget increases, the ROC-AUC and AP score experiences a significant drop.
@@ 5.5.2 Evaluation of untargeted attack
Table 3: Comparison of statistics of the two graphs.
5.6 Graph Visualization
@@
Upon observing Fig. 6 (a) and 6 (b), it is apparent that the two graph visualizations look very similar. Most of the communities are almost identical, indicating that the local structure of the graph is preserved. Since the ratio of modified links is much smaller than the total number of links, it is challenging to discern the detailed adversarial link modifications resulting from the proposed poisoned attack. These changes can be considered unnoticeable in the original graph. However, when observing Fig. 6 (c) and 6 (d), we notice that the overall appearance appears somewhat different. Most of the communities have changed, and even the sizes of the communities have changed. Upon closer inspection, Figs. 6 (c) and 6 (d) are more interconnected than Figs. 6 (a) and 6 (b). In addition, some new links have appeared in the upper right corner. To further investigate, we calculated the statistics of these graphs, as shown in Table 3.
From Table 3, we observe that the statistics of Figs. 6 (a) and 6 (b) are very similar, only have slight differences in some aspects. The statistic of Fig. 6 (c) and 6 (d) have much more difference, especially on Fig. 6 (d). We also conclude some critical findings from the visualization and statistic table. Firstly, we noticed that our adversarial modifications add more links to the original graph. Secondly, the average degree, average clustering coefficient, and average path length also indicate that the poisoned graph is more interconnected than the original graph. As the attack budget increases, the graph becomes more interconnected. These observations collectively demonstrate that the proposed attack slightly enhances the connectivity of the original graph. This enhancement is likely attributable to the fact that link adversarial activation performs better than link adversarial dissolution in targeted attacks.
6 Conclusion and Future work
In this paper, we introduce an innovative unweighted graph poisoning attack leveraging meta-learning techniques to undermine the link prediction performance of VGAE. Our approach capitalizes on meta-learning to identify optimal attack directions, ensuring subtle modifications in the graph. Through extensive experimental evaluation, we showcased the efficacy of our attack in targeted and untargeted scenarios. Our results demonstrate superiority over existing state-of-the-art methods.
For future research, we plan to explore three potential directions:
  • Improving link adversarial dissolution: Our approach exhibits limited effectiveness in breaking existing links. Future work could focus on improving the performance of this aspect of the attack.
  • Application in recommendation systems: Investigate the applicability of our approach in real-world recommendation systems, which often rely on user-item graphs for personalized recommendations.
  • Extension to weighted graph attack: Extending our current work to address graph adversarial attacks on weighted graphs expands the scope of our approach to handle more complex graph structures.

### v2

@@ 6   Compute attack loss $\widetilde{L}_{i}$;
7   Select $w_{i}$ from $W$;
8 $\widetilde{L}=\widetilde{L}+w_{i}\times\widetilde{L}_{i}$;
@@ 9   Back propagation on $\nabla$: update $\theta_{i}\rightarrow\theta_{i+1}$;
10
11
12Compute the meta-gradient $\widetilde{\nabla}$ with respect to attack target based on attack loss ($\widetilde{L}$);
13
[…]
15
16return $A_{adv}$;
@@ 1 Initial $A_{adv}=A$
2 Symmetry the gradient matrix $\widetilde{\nabla}$ = ($\widetilde{\nabla}$ + $\widetilde{\nabla}^{T}$)/2
@@
[…]
5 if $\widetilde{\nabla}_{ij}>0$ and $A_{ij}=0$ then
@@ 7    $\Delta$ = $\Delta$ - 1;
9 if $\widetilde{\nabla}_{ij}<0$ and $A_{ij}=1$ then
@@ IV-D Attack Strategy
IV-D1 Attack Model
The primary purpose of the attack model is to extract information by mimicking the original process since the attacker lacks access to the original model and the training process. The essence of the attack model (Algorithm 1) lies in leveraging weighted meta-learning during its training phase, enabling the attacker to gain a comprehensive understanding of the process. The attack model is trained on the original graph data using weighted cross-entropy (Algorithm 1, lines 3 and 4), defined as:
@@ IV-D1 Attack Model
The attacker leveraging the meta-learning idea to aggregate the attack loss during each epoch (Algorithm 1, line 6). The key part of the attack model is that the attacker considers each training epoch as a task (episode) and intends to learn from all tasks (epoches). Instead of looking into a specific phase in the training process, meta-learning helps the attacker to capture the overall information of the whole training. The attacker tries to maximize the attack loss with respect to the graph structure ($Adj$) to conduct the effective attack on the graph structure. At the end of the training, the attacker computes the meta-gradient from the attack loss for further processing (Algorithm 1, line 8).
IV-D2 Weighted Scheme
Meta-gradient is constructed to guide the subsequent adversarial modifications. Hence, it is critical to form an informative and effective meta-gradient. To enhance its effectiveness, we propose three distinct weighting schemes that assign varying importance to gradient information from each training epoch (Algorithm 1, line 7). These schemes determine how each epoch’s gradient contributes to the final aggregated meta-gradient:
  • Magnitude-based Weighting:
    Magnitude-based weighting builds on fundamental optimization principles and relates to influence functions  [25]. The gradient’s magnitude directly measures the loss function’s sensitivity to input changes. This scheme assumes that larger gradients indicate epochs of greater sensitivity to structural perturbations and thus higher importance. Accordingly, the gradient in epoch $i$ is weighted by its L2 norm to capture this effect:
    $w_{i}=\left\|\widetilde{\nabla}\tilde{L}_{i}\right\|_{2}$ (6)
  • Performance-based Weighting:
    The core idea of boosting algorithms [26] like AdaBoost indicates that the final prediction is a weighted combination of models, with more accurate models receiving a greater say. Similarly, our performance-based scheme uses the attack model’s own performance as a signal for attack reliability. This scheme links the epoch importance directly to the empirical performance of the attack model. The weight for each epoch is related to its Average Precision (AP) score on the validation data of link prediction task. It assumes that gradient is more informative when the model has a good performance (understanding) of the validation data. The weight for epoch $i$ is calculated as:
    $w_{i}=AP_{i}$ (7)
  • Linear-based Weighting:
    Inspired by curriculum learning [27], where models learn more effectively when data is introduced from simple to complex, our linear weighting adopts a similar philosophy. We hypothesize that gradients from later epochs carry greater attack-relevant information. Thus, weights increase linearly with training progress. As the model captures the basic data structure in early epochs, this scheme prioritizes gradients from later stages. The weight for epoch $i$ is normalized by the total number of epochs $k$ and computed as:
    $w_{i}=i/k$ (8)
IV-D3 Adversarial modification
@@
V Experimental Results
@@
V-A Experimental Setup
To evaluate the effectiveness of our proposed attack method, we implemented the entire architecture using PyTorch. Our experiments are designed to answer the following questions:
  • Does our approach effectively reduce link prediction performance?
  • What is the most effective weighted scheme?
@@ V-A Experimental Setup
To answer these questions, we deployed VGAE as the attack surrogate model and evaluated the effectiveness of our poisoning attack approach in various graph-based use cases, including citation network (Cora), social network (Twitch) and recommendation network (MovieLens). Each use case is evaluated on a state-of-the-art link prediction model with the corresponding dataset. By training link prediction models separately on the original and poisoned graph data, we can observe any disparities in link prediction performance, indicating the effectiveness of our attack. In addition, we compared the effectiveness of our approach with other state-of-the-art methods, namely PGD [13], DICE [14], IGA [15], CLGA [18] and MetaBase [17]. The implementations for PGD, DICE, and MetaBase were adapted from the Graph Reliability Toolbox [28], and CLGA was sourced from the authors’ public repository. As an official implementation for IGA was unavailable, we implemented it ourselves based on the description in the original paper.
All experiments were conducted on a server equipped with an NVIDIA 3090 GPU. We utilize a 2 layer VGAE (with 32 and 16 hidden units) as the attack surrogate model to generate adversarial modifications to the graph structure. The effectiveness and transferability of these attacks are then evaluated against three link prediction victim models: 1) a two-layer Graph Attention Network (GAT) with 32 and 16 hidden units, 2) a two-layer VGAE with 32 and 16 hidden units, and 3) a three-layer LightGCN with a 32-dimensional embedding space. As described in the previous section, the adversarial modification is constrained by the attack budget, which we treat as a key experimental parameter. To keep the attacks inconspicuous, we limit the budget to a small range, between 1% and 5% of the total links in the dataset, when generating poisoned graphs. For each experiment, we train both the original and the poisoned models using their respective datasets and evaluate them on the same test set to assess link prediction performance.
V-B Experimental Dataset and Models
Since the proposed approach requires to access the full adjacency matrix to gain a global perspective of the graph structure to guide the attack most effectively, the primary bottleneck is its space complexity ($O(N^{2})$). This constraint creates a trade-off between our attack’s global effectiveness and its scalability. To prioritize a rigorous validation of the novel attack algorithm itself, we selected three widely used graph datasets in the literature to evaluate the performance of our proposed attack approach. These datasets represent the use cases in citation network, social network, and recommendation network.
@@ V-B Experimental Dataset and Models
  • Cora (citation network): The Cora dataset [29] contains scientific publications that are categorized into 7 classes. It has 2,708 nodes and 5,429 links. Each node is described by a word vector that indicates the presence of certain words in the publication from a predefined dictionary. This dataset is used to evaluate the GAT model.
  • Twitch (social network): The Twitch Gamers dataset [30] is a social network of users from the Twitch streaming platform. The version for English-speaking users contains 7,126 nodes and 35,324 links, where nodes are users and links represent mutual friendships. Node features are created based on user profiles and streaming activities. This dataset is used to evaluate the VGAE model.
  • MovieLens (recommendation network): The MovieLens dataset [31] consisting of 600 users and 9,000 movies. The graph is represented as a bipartite graph, where edges connect users to the movies they have rated. The features are represented by learnable embeddings to evaluate the LightGCN model.
@@
(b) Twitch dataset
(c) MovieLens dataset
Figure 3: Comparsion using ROC-AUC and NDCG@20 metrics
@@
(b) Twitch dataset
(c) MovieLens dataset
Figure 4: Comparsion using AP and Recall metrics
V-C Evaluation Metric
@@
  • ROC-AUC and AP Score: For citation and social networks, we use the ROC-AUC (Area Under the ROC Curve) and the Average Precision (AP) score to evaluate the performance of the model after the poisoning attack. ROC-AUC score provides a comprehensive understanding of the model’s ability to distinguish between positive and negative links. It evaluates the performance of the model across all possible thresholds and is more resistant to class imbalance compared to accuracy, making it a robust measure. As graph data are usually considered imbalance data, the AP score emphasizes the precision-recall trade-off of the model. By comparing clean model training with poisoning training performance, the effectiveness of the poisoning attack can be clearly demonstrated, highlighting its impact on the robustness and performance of the model.
  • NDCG@k and recall: For the recommendation network, the evaluation focuses on the quality of the top-K item ranking rather than the simple prediction of links. We employ Normalized Discounted Cumulative Gain (NDCG@k) and Recall for evaluation. NDCG@k evaluates the quality of the ranking itself. It is a position-aware metric that assigns greater importance to relevant items placed higher on the list, making it highly sensitive to recommendation order. Recall measures the fraction of relevant items from the test set that are successfully included in the top K recommended list. A significant drop in these top K ranking metrics after our poisoning attack demonstrates the attack’s effectiveness in degrading the model’s ability.
TABLE I: Performance drop on the Cora dataset
Metric Attack Budget Uniform Performance Magnitude Linear
[…]
2.5% 0.0388 0.0401 0.0326 0.0403
5% 0.0526 0.0570 0.0487 0.0573
[…]
2.5% 0.0426 0.0441 0.0359 0.0444
5% 0.0562 0.0620 0.0528 0.0620
TABLE II: Performance drop on the Twitch dataset
Metric Attack Budget Uniform Performance Magnitude Linear
[…]
2.5% 0.0353 0.0416 0.0326 0.0487
5% 0.0504 0.0552 0.0456 0.0804
[…]
2.5% 0.0417 0.0426 0.0405 0.0523
5% 0.0508 0.0567 0.0477 0.0859
TABLE III: Performance drop on the MovieLens dataset
Metric Attack Budget Uniform Performance Magnitude Linear
[…]
2.5% 0.0156 0.0169 0.0073 0.0186
5% 0.0210 0.0171 0.0154 0.0295
[…]
2.5% 0.0198 0.0187 0.0092 0.0239
5% 0.0219 0.0231 0.0123 0.0328
@@
TABLE IV: Comparison of statistics of the two graphs.
V-D Effectiveness Analysis
V-D1 Evaluation of weighted scheme
[…]
Among the three proposed schemes, the linear-based scheme consistently emerges as the most impactful and powerful weighted scheme. It achieves the highest performance drop in the vast majority of experimental settings, showcasing its robustness across different data domains and evaluation metrics. The performance-based scheme also proves to be highly effective, delivering results that are often comparable to the linear scheme. In contrast, the magnitude-based scheme, while still superior to the uniform baseline, generally yields the smallest performance drop of our three proposed methods. The findings confirm that attack guided by the proposed weighting schemes can compromise far more effectively than an unguided, uniform attack, especially those using linear-based and performance-based weighting schemes.
V-D2 Evaluation of poisoning attack
To evaluate the effectiveness of the proposed poisoning attack, we benchmarked it using the most impactful weighted scheme (linear-based) against several state-of-the-art adversarial attack baselines, shown in Fig. 3 and Fig. 4. In the experiments, we used two baselines, the original model and the null model [32]. The original model represents the performance of the link prediction on the original graph data. It is unaffected by the attack budget as it is clean data. The effectiveness of the attack methods is evaluated by observing the drop on the evaluation metrics. The null model serves as a reference point to evaluate whether the approach is capable of highlighting the vulnerabilities of the graph data. All approaches demonstrate their effectiveness compared to the original model and null model. As shown in the figures, the proposed approach consistently achieves a significant drop in the victim model’s performance, slightly surpassing other state-of-the-art approaches. As the attack budget increases, the performance of the models experiences a significant drop. In particular, our method achieves a lower performance score while requiring a smaller attack budget. In conclusion, the comparison provides evidence for the superiority of our proposed method, demonstrating its advanced capability to identify and exploit critical vulnerabilities in link prediction models.
V-E Graph Visualization
@@
Upon observing Fig. 5 (a) and 5 (b), it is apparent that the two graph visualizations look very similar. Most of the communities are almost identical, indicating that the local structure of the graph is preserved. Since the ratio of modified links is much smaller than the total number of links, it is challenging to discern the detailed adversarial link modifications resulting from the proposed poisoned attack. These changes can be considered unnoticeable in the original graph. However, when observing Fig. 5 (c) and 5 (d), we notice that the overall appearance appears somewhat different. Most of the communities have changed, and even the sizes of the communities have changed. Upon closer inspection, Figs. 5 (c) and 5 (d) are more interconnected than Figs. 5 (a) and 5 (b). In addition, some new links have appeared in the upper right corner. To further investigate, we calculated the statistics of these graphs, as shown in Table IV.
From Table IV, we observe that the statistics of Figs. 5 (a) and 5 (b) are very similar, only have slight differences in some aspects. The statistic of Fig. 5 (c) and 5 (d) have much more difference, especially on Fig. 5 (d). From the visualization and statistical analysis, we draw several key observations. First, our adversarial modifications tend to add links to the original graph. Second, metrics such as average degree, average clustering coefficient, and average path length all show that the poisoned graph is more interconnected than the original. As the attack budget increases, this connectivity further intensifies.
VI Conclusion
In this article, we present an innovative unweighted graph poisoning attack that leverages meta-learning to degrade link-prediction performance. Our method employs a meta-learning framework with weighted schemes to identify optimal attack directions, enabling subtle yet effective modifications to the graph. Extensive experiments demonstrate the superior effectiveness of our attack and show the capability to identify the vulnerabilities of link prediction models.